Roy decides to buy a personal residence and goes to the bank for a $150,000 loan. The bank tells him that he can borrow the funds at 4% if his father will guarantee the debt. Roy's father, Hal, owns a $150,000 CD currently yielding 3.5%. The Federal rate is 3%. Hal agrees to either of the following:

  • Roy borrows from the bank with Hal's guarantee to the bank.
  • Cash in the CD (with no penalty), and lend Roy the funds at 2% interest.

Hal is in the 32% marginal tax bracket. Roy, whose only source of income is his salary, is in the 12% marginal tax bracket. The interest Roy pays on the mortgage will be deductible by him.

Considering only the tax consequences, answer the following.

a. The loan guarantee:
Hal's interest income from the CDs would be $ before taxes and $ after taxes.

Roy's interest expense from the bank loan would be $ before taxes and $ after taxes.

This arrangement would produce an overall negative  cash flow after taxes to the family of $.

b. The loan from Hal to Roy:
Hal's tax on the imputed interest income from the loan to Roy would be $.

Roy's tax benefit from the imputed interest expense from Hal's loan would be $.

This arrangement would produce an overall negative  cash flow after taxes to the family of $.

c. Which option will maximize the family's after-tax wealth?
Answer #1

a. The Loan Guarantee arrangement

For Hal's Interest Income from Current Deposits

Before Tax After Taxes

($150000*3.5%)=$5250 ($150000*3.5%)-($150000*3.5%*32%)-(5250*3%)=3728

For Roy's Interest Expenses after Taxes

Before Tax After Taxes

($150000*4%)= $6000 ($150000*4%)-($150000*4%*12%)= 5280

Net Negative Outflow of Family after Taxes

5280- 3728= $1553

Note : Assuming the Federal Tax paid on Interest Income is available for deduction.

($5250*3

b. Loan From Hal to Roy

For Hal's Tax on Imputed Interest from Roy

($150000*2%)- ($150000*2%*32%)= 2040

Loss of Interest from Current Deposits = $3728

Outflow is $1688

For Roy's Interest on Loan from Hal

Interest Outflow = ($150000*2)- ($150000*2%*12)= $2640

Family Outflow =$2640+$1688= $4328

Note: The interest income Hal would have earned from Current deposit is also considered as his deemed loss.

3. First Case of Bank Guarantee arrangement is considered suitable as it would generate lower outflow for family